OpenStudy (anonymous):

Solve the equation. Give an exact solution. log4(x-6)+log4(x-6)=1 A)8 B)37 C)-37,37 D)-8,8

OpenStudy (anonymous):

c is square root -37 and 37

OpenStudy (loser66):

You mean \(log_4(x-6)\) right? I mean base of log is 4, right?

OpenStudy (anonymous):

yea

OpenStudy (loser66):

I got A)

OpenStudy (anonymous):

how is the process though? i dont know what formula to use

OpenStudy (loser66):

Since you have options, then use it. Try x =8, then \(log_4( 8-6) = log_4(2)\) so the left hand side is \(log_4(2) + log_4(2) = 2log_4(2) = log_4(2^2)=log_44=1\)

OpenStudy (loser66):

If you need formula, then use this: \(log_a b + log_a c= log_a (b*c)\)

OpenStudy (anonymous):

much obliged

OpenStudy (loser66):

Since your 2 terms are the same, (the formula still hold)

OpenStudy (loser66):

you just apply it.

OpenStudy (loser66):

another way by using the convenience of the same term: \(log_4(x-6) + log_4(x-6) =2log_4(x-6) =log_4(x-6)^2=1\) and you know that \(1= log_44\) hence \(log_4(x-6)^2=log_44\), that allows you conclude that \((x-6)^2= 4\\x-6=\pm2\), hence x = 8 or x =4 . We reject x =4 because it makes \(log_4(x-6) = log_4(-2)\) which is undefined.

OpenStudy (loser66):

So far, I gave you 3 way to get the answer, right? hehehe..

OpenStudy (triciaal):